OpenStudy (anonymous):

Find the perimeter of the rectangle. A square with the sides x + 3 and 2x minus 4. A. 3x - 1 B. 6x - 1 C. 6x - 2 D. 4x - 1 I got A , correct ?

OpenStudy (anonymous):

Remember, P = 2W + 2L. L = x + 3 W = 2x - 4 Substitute and simplify.

OpenStudy (anonymous):

i tried that

OpenStudy (anonymous):

\[2(x+3) + 2(2x - 4) = (2x + 6) +(4x-8) = (2x+4x) + (6 -8) = 6x - 2\]

OpenStudy (anonymous):

The answer is C

OpenStudy (swag):

c